OpenStudy (anonymous):

If the height of a right circular cone is decreased by 8%, by what percent must the radius of the base be decreased so that the volume of the cone is decreased by 15%?

OpenStudy (amistre64):

Vcone = (1/3)(base area)(h)

OpenStudy (amistre64):

kinda depends on the base area to begin with I think

OpenStudy (anonymous):

Problem doesn't provide any measurements at all :P

OpenStudy (amistre64):

\[V=\frac{\pi. r^2.h}{3}\]

OpenStudy (amistre64):

\[V(1-.15)=\frac{\pi.r^2.h(1-.08)}{3}\]it looks like

OpenStudy (amistre64):

\[r^2*(\frac{1}{.85})\]looks to be what they expet you to work with then

OpenStudy (anonymous):

The given answer is 4%.

OpenStudy (amistre64):

how to work that in i dunno

OpenStudy (amistre64):

sqrt(r^2sqrt(1/.85)) maybe

OpenStudy (amistre64):

r(1.04) is what the comes to; but that seems to increase r

OpenStudy (amistre64):

.85/.92 maybe? nope... lol

OpenStudy (anonymous):

x^2=.85/.92

OpenStudy (anonymous):

that is what you multiply the radius by to decrease volume 15 % you get .96 which is a 4 % decrease

OpenStudy (amistre64):

i was thiiiisssss lose lol

OpenStudy (anonymous):

\[\frac{1}{3}.92h \pi(xr)^2=.85 \frac{1}{3}\pi r^2h\]

OpenStudy (anonymous):

the x gets squared, most of the other stuff drops out